Telescope Bootstrap Startup: 14 September 2012 UT

Observer: DMiller (LBTO), OKuhn, MEdwards(Tucson)
Telescope Operator: GBechetti (LBTO)
Software Support: CBiddick (Tucson)
AO Support: DMiller (Tucson)
Telescope Support: MMidkiff (LBTO)
Instrument Support: ESolheid (LBTO)

Summary

There were scattered clouds and fog all night, and the humidity never dropped below 98%

Olga and Michelle performed some tests and calibrations with MODS1 and the LBCs.

Vidhya and Alfio also ran some LBTI TCS/AOS test.

Closed Dome Tests

04:30 Still cloudy up at the telescope. Retroreflector removed and fills done.

04:30 Waking MODS up for another round of calibrations. Dome dark, telescope at zenith.

04:47 sieveSnap (mods1[b,r].20120914.0001.fits), followed by
  • gg495.cal script.
    • r images 2,3 are from gg495 (red grating with GG495 (2) and clear (3))
  • wiggle.cal script
    • b 2-6 and r 408

05:17 done with wiggle.cal. Will take pixel flats again: grpixflats (24 minutes), grpixflats_r (14 min) and grpixflats_b (23 min)
  • in grpixflats_b --- look at flat with UG5 (b 0023)--- there is a bright circular arc neat bottom center. Well, I looked back at some blue-only UG5 pixflats from 20120212 (87), but it used VFLAT 10 and was very saturated at the red end, so any gradients near the center are hard to see, but I think I do see a hint of this circular arc there, too. I don't see this in the Blue-only Clear pixflats. And in the Dual UG5 flats, it looks like the brightest region coincides with the position of this arc.

05:31 mods data from tonight are being copied to /newdata fine, but they are going into /Repository/20120913 (yesterday's) not /Repository/20120914/ IT4218

06:26 start grflats.cal (to 06:59, 33 min); grflat_r (13 min) and grflat_b (lost track but I think 23 min)

06:30 Geno authorizes LBC on both sides. Vidhya was doing some testing with LBTI - but done now.

07:15 Tried lbcrangebal script after the modifications for 1/ a sanity check and 2/ logging to /home/lbceng/RBlogs/ I also edited the portion which copies LBCCHIP2 extension image.fits[2] to img because when the script makes a copy of an image, it takes a very long time --- this was a limitation in the old lbcrangebal.
Works OK --- didn't display the image to RB_Science ds9 - had to open a separate ds9 - but after that, it displayed the images, albeit only a portion of chip2 ( Y from 257 to 4352 and X from 2 to 3200)... imt issue? or because I used a 4-chip image?

07:23 Running FILTER testing OB to cycle through filters.

07:26 Lights needed for filter checks - looks OK, though through some filters not much light (too short exposures)... oh, well lbc is not in the beam so it's

07:55 Lights off and sending a test preset from LBC which will take a pair of 60-sec exposures and dither 3 times.
  • btec chip levels are not equalized btec_1 has background ~100 counts while btec_2 has background ~600 counts
  • rtec chip backgrounds are ~300.
  • the preset completes with warning but that is only because the swing arms are not configured correctly for LBC:
Fri Sep 14 08:04:47.158 2012 Preset command wait for telescope complete for both LBC
Fri Sep 14 08:04:47.217 2012 Preset alert for both LBC mode TRACK: Left side swing arm configuration wrong
Fri Sep 14 08:04:47.322 2012 Preset alert for both LBC mode TRACK: Right side swing arm configuration wrong
Fri Sep 14 08:04:47.372 2012 Preset alert for both LBC mode TRACK: Stop[22] complete
Fri Sep 14 08:04:47.423 2012 Preset alert for both LBC mode TRACK: Stop[9] complete
Fri Sep 14 08:04:47.473 2012 Preset command complete with warnings for both LBC mode TRACK with offset, Sound=presetWarning.wav

08:24 grlamps --- done at 08:53. The blue-only argon (last one mods1b.20120914.0051.fits) has a reflection of the slit at X~3127 --- haven't noticed this, but haven't looked in that much detail before.

09:12 doBias --- had to run red fitsflush to get 52,53,54. MODS going to sleep after another night of calibrations.

09:20 RH is still 101% ... doesn't look very promising. We are all, in Tucson and LBT, calling it a night.

-- OlgaKuhn - 14 Sep 2012
Topic revision: r3 - 14 Sep 2012, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback